Y/4+5=2(y-1)
I having a hard time understanding how to solve this problem.

Answer by Alan3354(69443)

y/4+5=2(y-1)
y/4 + 5 = 2y - 2
Multiply thru by 4
y + 20 = 8y - 16
Subtract y
20 = 7y - 16
Add 16
36 = 7y
Divide by 7
y = 36/7